Circular No. 5636

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SWIFT-TUTTLE (1992t) Orbital computations by the undersigned, and also by S. Nakano,

Sumoto, Japan, have so far failed to link all the observations, even when allowance is made for nongravitational forces. Although a reasonable fit can be made to the 1862 (except October) and 1992 observations, the resulting transverse nongravitational component is so large that the resulting eighteenth-century perihelion time is 15 months too late. Alternatively, although the three perihelion times can be well represented without any consideration of nongravitational forces at all, there are strong systematic errors, amounting to more than 1′, in 1862 and 1992. The gravitational orbital elements below satisfy the observations in 1992 and in Oct. 1862 very well, and they also represent the presumed 1737 perihelion time within 1 day. Backward computation of this solution reveals few candidates for earlier appearances of the comet, although the one of –68 fits within 1 year (there being 15 revolutions between then and 1862), and the comet of +60 may also belong. Future extrapolation gives the next return to perihelion as 2126 July 11, although the problem with the computation of the nongravitational forces must introduce some uncertainty; a change by +15 days could cause the comet to hit the earth on 2126 Aug. 14. It therefore seems prudent to attempt to follow P/Swift-Tuttle for as long as possible after the present perihelion passage, in the hope that an adequate independent orbit determination, uncontaminated by nongravitational effects, can be made from mid-1993 (at r = 3 AU and far to the south) to, say, 1998 (when r = 15 AU and an assumed nuclear absolute magnitude of 14 yields an apparent magnitude of 26).

Epoch = 1992 Dec. 4.0 TT

T = 1992 Dec. 12.323 TT  ω = 153.^{o}013
e = 0.96359  Ω = 139.456 }2000.0
q = 0.95812 AU  i = 113.430

a = 26.31666 AU n^{o} = 0.007301 P = 135.00 years Total visual magnitude estimates: Oct. 13.01 UT, 7.9 (H. Lüthen,

Hamburg, Germany, 0.10-m reflector); 13.76, 7.6 (S. Canonaco, Genk, Belgium, 0.20-m refractor); 14.74, 7.8 (H. Mikuz, Ljubljana, Slovenia, 7 × 50 binoculars); 14.86, 7.6 (M. Tombelli, Montelupo, Italy, 0.20-m reflector).

1992 October 15 Brian G. Marsden